OpenStudy (anonymous):

What is the antiderivative of x^1/3?

OpenStudy (anonymous):

use the power rule backward \[\int x^ndx=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

something missing there satellite

OpenStudy (anonymous):

yeah the stupid and quite irrelevant \(+C\)

OpenStudy (anonymous):

:p

OpenStudy (anonymous):

not that irrelevant. It is very important (at least in engineering when you're calculating a bending beam)

OpenStudy (anonymous):

Sorry, I missed the messages on this side of the screen. I'm calculating the average value of g(x) = x^1/3 on [1,8]. I'm drawing a blank on integrating a fraction...

OpenStudy (anonymous):

(1/3) = n just apply satellite's formula

OpenStudy (anonymous):

1/2x^4/3 ?

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }x ^{\frac{ 4 }{ 3 }}\]

OpenStudy (anonymous):

I don't understand where the 3/4 comes from...?

OpenStudy (anonymous):

\[\frac{ x ^{\frac{ 4 }{3 }} }{ \frac{ 1 }{ 3}+1 }\]

OpenStudy (anonymous):

Got it! Thanks!